OpenStudy (anonymous):

What is the exact value of the expression the square root of 4 times y to the eighth power. + 6the square root of y squared. − 2the square root of y squared. + the square root of y to the eighth power.? Simplify if possible. A:5y4 + 3y B:7y4 C:3y4 + 4y D:7y6

OpenStudy (anonymous):

yes

OpenStudy (ganpat):

i guess, i wrote something wrong.. well this is an answer.. i will correct myself.. \[2y ^{4} + 6y - 2y + y ^{4} = 3y ^{y} + 4y\] option C !!

OpenStudy (ganpat):

\[\sqrt{4(y ^{8})} + 6\sqrt{y ^{2}} - 2 \sqrt{y ^{2}} + \sqrt{y ^{8}}\] this was the right equation..

OpenStudy (radar):

\[2y ^{4}+6y-2y+y ^{4}\]Combine like terms. \[3y ^{4}+4y\]

OpenStudy (anonymous):

7y^4

OpenStudy (radar):

You should not attempt to combine unlike terms. \[y ^{4}\]is not similar to y.

OpenStudy (ganpat):

nopes.. terms with same powers are to be added.. with different powers cannot be !!